Two angles are complementry.The smaller angle is 15 degrees more than half the larger angles.What is the measure of the larger angle?
Answer by laoman(51)

If two angles are complementary, they add up to 90°
If we know that, let's assume the smaller angle is x and the larger angle is y
From the first statement, x+y = 90 -----eqn1
From the second statement, x = 15 + y/2 ---eqn2
Solving the two eqns simultaneously,
i.e From eqn1, x = 90-y after making y the subject of formula
substitute x into eqn2

90-y = 15+y/2
Let's multiply althrough by 2
180-2*y = 30+y
Let's take like terms and solve for y
3*y = 180-30
y=150/3 = 50°
Therefore the larger angle (which is y) is 50°
QED